OpenStudy (anonymous):

Michael works two part time jobs for a total of 23 hours a week. Job A pays 6.20 per hour and job B pay 6.60 per hour. How many hours did he work at each job the week that he made 148.20?

OpenStudy (anonymous):

assume he worked x hours at Job A, y hours at Job B 6.2x + 6.6y = 148.2 x + y = 23

OpenStudy (anonymous):

how many hours for job a and b?

OpenStudy (anonymous):

6.2x + 6.2y = 6.2*23 6.2x + 6.6y = 148.2 0.4y = 148.2 - 6.2*23 y = 14 so x = 9

OpenStudy (anonymous):

Thanks you so much!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

If Michael wirks t hours at job An then he works 23 - t houre at job B. Total earnings for job A is 6.2t dollars, and for job B it's 6.6(23-t) dollars. In terms of t, his total from both jobs that week is 6.2t + 6.6(23-t) = 151.8 - .4t = 148.2. .4t = 3.6 t = 36/4 = 9. He worked 9 hours at job A and 14 hours at job b.

OpenStudy (anonymous):

oh thank you so much!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

:^)